OpenStudy (anonymous):

Rewrite the equation below in standard form. -3x+6y=12 A. 3x-6y=12 B. 3x-6y=-12 C. -x+2y=4 D. x-2y=-4

OpenStudy (anonymous):

why not b. though?

OpenStudy (texaschic101):

standard form is Ax + By = C -3x + 6y = 12 (we do not want x to be negative, so we will multiply the entire equation by -1) -1(3x + 6y) = -1(12) 3x - 6y = -12 --> standard form but I think you can divide by 3 and the answer would be : x - 2y = -4 can there be more then one answer ?

OpenStudy (texaschic101):

I would go with B.

OpenStudy (anonymous):

ok